# Group-robust Machine Unlearning

### v1

@@ 4.3 Impact of different unlearning ratios
Figure 5 analyzes methods Avg. Gap across different unlearning ratios. On CelebA [31], all methods show similar discrepancies from Retrain + reweight at a low unlearning ratio (i.e., 0.1), likely because the retain set’s group statistics are close to the training distribution. Controversy, all algorithms struggle at 0.1 unlearning ratio in Waterbirds [42], where the small forget set size causes UA fluctuations and affects the BatchNorm [22] estimation. As the unlearning ratio grows (i.e., 0.5 and 0.9), MIU outperforms baselines by a growing margin. At unlearning ratio 0.9, MIU achieves 98.3% on CelebA [31] (vs. 94.8% of L1-sparse [23]), and 97.2% on Waterbirds [42] (vs. 93.6% of SCRUB [28]). Furthermore, MIU remains more consistent across unlearning ratios, confirming that our design choices effectively narrow the Avg. Gap with Retrain + reweight. Full Fig. 5 tables are reported in the Appendix.
@@ 4.4 Multi-group unlearning
This section compares MIU and existing approaches in multi-group unlearning, i.e., when the forget set data is sampled from multiple groups. Figure 6 shows the outcome of this experiment on FairFace [24], given its numerous groups, with a varying number of groups in the forget set and an unlearning ratio fixed to 0.5. All methods achieve high scores and a good discrepancy with Retrain + reweight, though SCRUB [28] performs worst (97.1% of Avg. Gap). Performance gaps shrink as more groups are included, as forget-set statistics align with the original training distribution, reducing non-uniform sampling effects. Nonetheless, MIU is the most consistent, highlighting its effectiveness in group-robust machine unlearning.
@@ 4.5 Ablations
| Dataset | Eq. 5 | Eq. 3 | Eq. 4 | RW | UA | GA | Gap |
| CelebA | 53.6 | 54.0 | 94.1 | ||||
| 55.3 | 55.9 | 94.9 | |||||
| 41.7 | 42.3 | 78.9 | |||||
| 68.8 | 69.0 | 99.2 | |||||
@@
Table 4 shows the ablation of MIU components in all three datasets by systematically adding each element to understand its contribution. We mark with “ ” when the component is used in the experiment and “” when it is not. From left to right, we list retaining term, unlearning term, calibration term, and reweight. In the first row, we consider our method when only the unlearning term and the retaining term are used. We highlight how the UA is low for all three datasets, demonstrating that mutual information minimization can be used to unlearn. This baseline already achieves remarkable Avg. Gap with Retrain + reweight, scoring a 94.1% in CelebA [31], which is the best among methods that do not use reweight. Adding our calibration term (Eq. 4), leads to an increase in GA in all three datasets, with FairFace [24] showing a growth of 6.9%. We highlight that the forget accuracy also grows. Yet, this increase is caused by Eq. 4 that calibrate the mutual information to match the original group robustness. Therefore, the high UA cannot be blamed on the poor unlearning. Although most previous approaches exploit a retaining term [9, 28, 23, 13], we also ablate this component for completeness. As previous works suggest [28, 9], removing the retaining term negatively impacts model utility, resulting in the lowest performance overall. Similarly, when adding reweight, the Avg. Gap gets improved in all three datasets, as we also highlight in Fig. 4. These results highlight the contribution of each of MIU components that allow for unlearning (Eq. 3) while preserving forget set dominant group performance (Eq. 4).
@@ 5 Conclusion
This paper is the first to address the performance degradation caused by non-uniformly distributed forget sets in both model retraining and approximate unlearning. We show that adopting a simple data distribution reweighting (reweight) for retraining the model is a simple and better alternative than retraining with group-DRO [42]. Moreover, we propose the first approximate unlearning method (MIU) that unlearns personal data while reducing the risk of degradation of the forget set dominant group. Our evaluation demonstrates that Retrain + reweight consistently improves over a simple Retrain while MIU outperforms existing baselines in group-robust machine unlearning.
@@
Acknowledgements. We acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support. E.R. and M.M. are supported by the MUR PNRR project FAIR - Future AI Research (PE00000013), funded by NextGeneration EU. E.R. is also supported by the EU projects AI4TRUST (No.101070190) and ELIAS (No.01120237) and the PRIN project LEGO-AI (Prot.2020TA3K9N). T.D.M. is funded by NextGeneration EU. This work has been supported by the French National Research Agency (ANR) with the ANR-20-CE23-0027. We thank Olivier Laurent for his valuable feedback and insightful suggestions.
Supplementary Material
| Method | reweight | RA | UA | TA | MIA | EO | GA | Avg. Gap |
| Pretrain | 83.80.0 | 67.90.5 | 82.70.1 | 0.10.1 | 20.81.0 | 69.00.9 | - | |
| Retrain | 83.80.1 | 62.01.6 | 82.50.1 | 0.20.1 | 21.91.3 | 64.01.9 | - | |
| Retrain | 83.70.2 | 63.21.6 | 82.70.2 | 0.20.0 | 21.20.5 | 65.81.7 | - | |
| L1-sparse [23] | 82.20.1 | 52.80.5 | 81.50.0 | 0.10.0 | 29.80.1 | 56.01.4 | 94.80.5 | |
| SalUn [13] | 83.10.8 | 69.28.5 | 81.80.9 | 0.10.1 | 23.81.8 | 69.18.0 | 96.92.0 | |
| SCRUB [28] | 84.40.0 | 64.90.6 | 82.90.2 | 0.10.0 | 22.00.3 | 65.00.1 | 99.10.2 | |
| MIU | 83.90.1 | 63.13.8 | 82.70.0 | 0.10.0 | 22.30.4 | 64.03.8 | 98.60.8 | |
| L1-sparse [23] | 82.20.2 | 60.72.0 | 81.50.2 | 0.10.0 | 27.40.6 | 63.42.2 | 97.30.8 | |
| SalUn [13] | 83.50.1 | 61.92.0 | 82.60.1 | 0.10.1 | 22.71.0 | 63.12.0 | 98.30.2 | |
| SCRUB [28] | 84.40.0 | 66.80.7 | 82.90.1 | 0.20.0 | 20.40.2 | 67.60.5 | 98.80.6 | |
| MIU | 83.80.2 | 67.25.6 | 82.40.2 | 0.10.1 | 20.90.6 | 67.84.3 | 98.81.2 |
| Method | reweight | RA | UA | TA | MIA | EO | GA | Avg. Gap |
| Pretrain | 84.60.1 | 66.61.8 | 82.80.0 | 0.10.0 | 20.90.7 | 67.32.4 | - | |
| Retrain | 84.70.4 | 54.64.1 | 82.30.1 | 0.10.1 | 31.60.9 | 56.94.3 | - | |
| Retrain | 84.50.1 | 66.62.4 | 82.40.1 | 0.30.1 | 20.80.4 | 67.81.6 | - | |
| L1-sparse [23] | 83.80.2 | 44.33.5 | 81.30.2 | 0.20.1 | 37.50.2 | 47.53.0 | 89.80.4 | |
| SalUn [13] | 84.90.2 | 47.64.1 | 82.40.2 | 0.10.1 | 31.60.7 | 48.83.9 | 91.70.8 | |
| SCRUB [28] | 82.12.3 | 40.64.8 | 79.82.4 | 0.40.1 | 40.12.7 | 42.65.4 | 87.43.6 | |
| MIU | 84.80.0 | 55.30.8 | 82.60.2 | 0.30.1 | 27.40.4 | 55.91.0 | 94.90.7 | |
| L1-sparse [23] | 83.40.1 | 60.61.8 | 81.30.2 | 0.20.0 | 27.90.4 | 62.02.1 | 96.51.2 | |
| SalUn [13] | 84.40.2 | 64.92.4 | 82.50.2 | 0.10.0 | 21.41.4 | 66.22.2 | 98.51.0 | |
| SCRUB [28] | 84.40.5 | 61.62.3 | 82.60.4 | 0.50.1 | 24.01.3 | 62.91.0 | 97.71.5 | |
| MIU | 84.20.1 | 68.80.3 | 82.50.1 | 0.10.0 | 20.20.6 | 69.01.2 | 99.20.6 |
| Method | reweight | RA | UA | TA | MIA | EO | GA | Avg. Gap |
| Pretrain | 85.30.1 | 67.60.9 | 82.70.1 | 0.10.1 | 20.00.8 | 68.00.9 | - | |
| Retrain | 86.80.4 | 27.82.4 | 80.70.2 | 1.30.2 | 50.01.0 | 29.92.2 | - | |
| Retrain | 84.20.9 | 64.24.5 | 81.80.2 | 0.40.4 | 22.90.9 | 66.84.8 | - | |
| L1-sparse [23] | 86.90.1 | 15.61.5 | 79.60.2 | 10.02.6 | 54.60.8 | 17.01.1 | 75.91.9 | |
| SalUn [13] | 87.00.3 | 17.310.0 | 80.01.0 | 8.25.4 | 44.40.8 | 18.39.9 | 78.42.9 | |
| SCRUB [28] | 71.51.2 | 34.01.6 | 67.01.1 | 0.90.9 | 36.63.3 | 34.62.1 | 82.61.1 | |
| MIU | 87.30.1 | 31.62.0 | 81.30.2 | 1.90.4 | 38.30.2 | 32.81.8 | 85.51.1 | |
| L1-sparse [23] | 85.00.4 | 56.03.2 | 80.60.2 | 5.12.1 | 31.10.4 | 58.52.7 | 94.81.5 | |
| SalUn [13] | 85.11.0 | 59.711.6 | 81.40.6 | 1.00.8 | 23.52.2 | 60.610.8 | 94.31.0 | |
| SCRUB [28] | 64.25.2 | 66.36.2 | 62.84.4 | 0.20.1 | 16.010.7 | 66.97.4 | 87.81.3 | |
| MIU | 82.93.1 | 64.13.7 | 79.82.1 | 1.51.2 | 25.83.1 | 66.44.6 | 98.30.6 |
| Method | reweight | RA | UA | TA | MIA | EO | GA | Avg. Gap |

### v2

@@ 4.3 Impact of different unlearning ratios
Figure˜6 analyzes methods avg. gap across different unlearning ratios. On CelebA (Liu et al., 2015), all methods show consistent discrepancies from Retrain + reweight across different unlearning ratios, likely because reweight recovers a large portion of the lost robustness. Nonetheless, MIU consistently outperforms existing approaches, scoring 99.4%, 99.0%, and 98.8% at unlearning ratios of 0.1, 0.5, and 0.9. Instead, all algorithms struggle at 0.1 unlearning ratio in Waterbirds (Sagawa et al., 2020), where the small forget set size (i.e., five samples) causes high fluctuations in the UA, as each sample counts as of the total classification error. Furthermore, the reduced forget set size makes the gradient and BatchNorm (Ioffe, 2015) estimation noisy. As the unlearning ratio grows (i.e., 0.5 and 0.9), MIU outperforms baselines by a growing margin. At unlearning ratio 0.9, MIU achieves 97.2% on Waterbirds (vs. 93.6% of SCRUB (Kurmanji et al., 2024)). Furthermore, MIU remains more consistent across unlearning ratios, confirming that our design choices effectively narrow the avg. gap with Retrain + reweight. Full Fig.˜6 tables are reported in Appx.˜B.
@@ 4.4 Multi-group unlearning
This section compares MIU and existing approaches in multi-group unlearning, i.e., when the forget set data is sampled from multiple groups (group composition is in Sec.˜B.3). Figure˜6 shows the outcome of this experiment on FairFace (Karkkainen & Joo, 2021), given its numerous groups, with a varying number of groups in the forget set and an unlearning ratio fixed to 0.5. All methods achieve high scores and a good discrepancy with Retrain + reweight, though SCRUB (Kurmanji et al., 2024) performs worst (97.1% of avg. gap). Performance gaps shrink as more groups are included, as forget-set statistics align with the original training distribution, reducing non-uniform sampling effects. Nonetheless, MIU is the most consistent, highlighting its effectiveness in group-robust machine unlearning.
@@ 4.5 Ablations
Table˜4 shows the ablation of MIU components in all three datasets by systematically adding each element to understand its contribution. We mark with “ ” when the component is used in the experiment and “” when it is not. From left to right, we list retaining term, unlearning term, calibration term, and reweight. In the first row, we consider our method when only the unlearning term and the retaining term are used. We highlight how the UA is low for all three datasets, demonstrating that mutual information minimization can be used to unlearn.
| dataset | Eq.˜7 | Eq.˜3 | Eq.˜6 | RW | UA | GA | gap |
| CelebA | 35.5 | 35.4 | 97.5 | ||||
| 36.3 | 36.3 | 98.0 | |||||
| 27.2 | 28.5 | 95.2 | |||||
| 43.2 | 41.3 | 99.0 | |||||
@@
This baseline already achieves a remarkable avg. gap with Retrain + reweight, scoring a 97.5% in CelebA (Liu et al., 2015), which is the best among methods that do not use reweight. Adding our calibration term (Eq.˜6), leads to an increase in GA in all three datasets, with FairFace (Karkkainen & Joo, 2021) showing a growth of 6.9%. We highlight that the forget accuracy also grows. Yet, this increase is caused by Eq.˜6, which calibrates the mutual information to match the original group robustness. Therefore, the high UA cannot be blamed on the poor unlearning. Although most previous approaches exploit a retaining term (Chundawat et al., 2023a; Kurmanji et al., 2024; Jia et al., 2023; Fan et al., 2024b), we also ablate this component for completeness. As previous works suggest (Kurmanji et al., 2024; Chundawat et al., 2023a), removing the retaining term negatively impacts model utility, resulting in the lowest performance overall (e.g., 81.9% in Waterbirds). Similarly, when adding reweight, the avg. gap gets improved in all three datasets (e.g., 96.9% in Waterbirds), as we also highlight in Fig.˜4. These results highlight the contribution of each of MIU components that allow for unlearning (Eq.˜3) while preserving forget set dominant group performance (Eq.˜6).
@@ 5 Conclusion
This paper is the first to address the performance degradation that can be caused by non-uniformly distributed forget sets in both model retraining and approximate unlearning. We show that adopting a simple data distribution reweighting (reweight) for retraining the model is a simple and better alternative than retraining with group-DRO (Sagawa et al., 2020). Moreover, we propose the first approximate unlearning method (MIU) that unlearns personal data while reducing the risk of degradation of the forget set dominant group. Our evaluation demonstrates that Retrain + reweight consistently improves over a simple Retrain while MIU outperforms existing baselines in group-robust machine unlearning.
Limitations. One limitation of our work is the assumption that group annotations are known, which may not hold in real-world applications where such labels are difficult to obtain. Therefore, a natural follow-up of this work is a group-agnostic methodology that preserves model robustness as achieved by reweight. A naïve approach towards this direction would be to discover groups from data (Kim et al., 2024; D’Incà et al., 2024), and applying approaches presented in this paper. Furthermore, our evaluation is restricted to the classification setting. Applying the proposed techniques and baselines to other domains might be non-trivial, and there is no guarantee that unlearning effectiveness and accuracy will transfer directly. Despite these limitations, our work is an important first step toward understanding and mitigating the accuracy degradation caused by group-unbalanced forget sets.
Broader impact statement. Machine unlearning is designed to remove user data from a trained model. While its primary goal is to preserve privacy, it can also be misused for malicious purposes. Targeted unlearning of specific groups may lead to biased models with harmful consequences. However, the proposed group-robust machine unlearning seeks to minimize performance degradation for dominant groups in the forget set. Thus, MIU and reweight counteract biases introduced by unlearning, mitigating the negative societal impacts of its misuse.
@@
Acknowledgements. We acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support. Elisa Ricci and Massimiliano Mancini are supported by the MUR PNRR project FAIR - Future AI Research (PE00000013), funded by NextGeneration EU. Elisa Ricci is also supported by the EU projects AI4TRUST (No.101070190) and ELIAS (No.01120237). Thomas De Min is funded by NextGeneration EU. This work is also supported by the EU project ELLIOT (101214398) and by the French National Research Agency (ANR) with the ANR-20-CE23-0027. We thank Olivier Laurent for his valuable feedback and insightful suggestions.
Appendix A Metrics
This section provides further details on the adopted metrics, i.e., the RA, UA, TA, the membership-inference attack (MIA) (Yeom et al., 2018), the dominant forget group accuracy (GA), the equalized odds (EO), and the average gap (avg. gap) (Fan et al., 2024b).
RA, UA, and TA. We evaluate the model on the remaining, forget, and test sets to compute the remaining, forget, and test accuracy. Therefore, we report the ratio of correctly classified samples for each of these subsets of the dataset.
MIA. To compute the membership inference attack, we follow previous works (Jia et al., 2023; Fan et al., 2024b) and train a model on remaining and validation losses to predict membership. Therefore, we assign a different binary label to remaining and validation losses and train the model to discriminate between them. After model training, we use such a model to infer the membership of forget set data. In all tables, we report the MIA-Efficacy (Jia et al., 2023) that is computed as follows:
where are the true negatives, i.e., the number of samples the MIA predicted as non-members. Instead of training a support vector machine as in (Jia et al., 2023; Fan et al., 2024b), we used a random forest as the accuracies are comparable with an SVM. Moreover, training is faster since it can be easily parallelized. The higher the MIA-Efficacy, the better the model’s privacy protection.
EO. Equalized Odds (Hardt et al., 2016) measures model fairness or prediction dependencies on protected attributes. To compute the EO, we measure model performance discrepancy by varying the sensitive attribute value and averaging over different target labels. Formally, the EO for a binary classification model is computed as follows (Hardt et al., 2016):
where are random variables describing model predictions, target attributes, and sensitive attributes. EO measures the absolute difference in outputting a positive prediction when the protected attribute equals 1 and 0, averaging over the two target attributes. In the FairFace (Karkkainen & Joo, 2021) case, we set all classes and protected attributes that do not match those of the dominant group of the forget set as and , as FairFace counts more than two classes and more than two attributes. The lower the EO, the better the model fairness.
avg. gap. Following previous works (Jia et al., 2023; Fan et al., 2024b), we compute the avg. gap to simplify the comparison among different methodologies. avg. gap is computed as the average metric discrepancy between the unlearned model and the retrained gold standard. Formally, let be the set of all metrics used in this paper, then avg. gap is computed as follows:
Appendix B Additional results
This section contains the additional results that could not be included in the main paper. Section˜B.1 intuitively shows the accuracy degradation caused by non-uniformly sampled forget sets in all three investigated datasets. Sections˜B.2 and B.3 report tables associated with experiments on different unlearning ratios and sampling the forget set from multiple groups. Section˜B.5 further expands MIU ablations. Finally, Section˜B.6 further evaluates MIU’s and reweight’s fairness and robustness preservation.
B.1 Unlearning non-uniformly sampled data
Figure˜2 of the main paper shows how the forget set dominant group accuracy drops for least represented groups when increasing the unlearning ratio, limiting the analysis to CelebA (Liu et al., 2015) due to space constraints. This section reports the same experiment on all three investigated datasets for completeness.
Figure˜7 shows accuracy variations when unlearning blond males (minority) and non-blonde females (majority) in CelebA (Liu et al., 2015), waterbirds on land (minority) and landbirds on land (majority) in Waterbirds (Sagawa et al., 2020), and more than 70 y.o. Middle Easterns (minority), 60-69 y.o. Caucasians (minority), 20-29 y.o. African Americans (majority), and 30-39 y.o. Southeast Asians (majority) in FairFace (Karkkainen & Joo, 2021), with different unlearning ratios: , , . While the accuracy drop is more evident on CelebA (Liu et al., 2015), it is also visible in the other two datasets. Interestingly, some unlearning methods are more robust to this performance drop than others (e.g., SCRUB (Kurmanji et al., 2024) scores a 56.25% in FairFace (Karkkainen & Joo, 2021) with unlearning ratio 0.9 vs. 38.54% of SalUn (Fan et al., 2024b) in minority groups, where the latter is the second best).
Although some methods are more robust, we argue that the quality of the unlearning process influences the accuracy of the dominant group of the forget set. The less the forget set was unlearned, the more the performance retention. As an example, Sec.˜4.2 highlights that likely SCRUB (Kurmanji et al., 2024) fails to effectively unlearn the forget set in the FairFace (Karkkainen & Joo, 2021) experiment, showing higher minority group accuracy (56.25%) than the Retrain (31.25%). Thus, to better investigate the relationship between unlearning effectiveness and performance retention, Sec.˜B.2 reports tables associated with Fig.˜7 using all investigated metrics.
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
| Pretrain | 96.20.0 | 44.02.7 | 95.90.1 | 0.50.7 | 23.40.5 | 41.91.7 | - | |
| Retrain | 96.20.0 | 39.91.0 | 95.80.0 | 0.20.3 | 24.60.3 | 39.60.7 | - | |
| Retrain | 96.20.0 | 44.42.1 | 95.90.0 | 0.70.6 | 23.40.7 | 42.42.2 | - | |
| L1-sparse (Jia et al., 2023) | 95.40.1 | 36.03.9 | 95.40.1 | 0.70.0 | 27.60.7 | 31.34.4 | 95.71.2 | |
| SalUn (Fan et al., 2024b) | 96.20.0 | 36.21.0 | 95.80.0 | 0.00.0 | 26.90.6 | 34.42.3 | 96.61.4 | |
| SCRUB (Kurmanji et al., 2024) | 96.40.0 | 42.82.1 | 96.00.0 | 0.50.7 | 24.50.3 | 40.60.8 | 98.80.8 | |
| MIU | 96.20.0 | 43.02.4 | 95.90.0 | 0.20.3 | 23.40.1 | 42.20.5 | 99.10.3 | |
| L1-sparse (Jia et al., 2023) | 95.40.0 | 38.91.9 | 95.40.0 | 1.00.3 | 26.31.4 | 36.54.1 | 97.31.0 | |
| SalUn (Fan et al., 2024b) | 96.10.0 | 40.15.9 | 95.90.0 | 0.70.6 | 25.01.9 | 38.76.1 | 97.20.9 | |
| SCRUB (Kurmanji et al., 2024) | 96.40.0 | 44.73.0 | 96.00.0 | 0.50.3 | 23.90.4 | 42.01.3 | 99.00.4 | |
| MIU | 96.20.0 | 44.92.7 | 95.90.0 | 0.20.3 | 22.90.5 | 43.00.7 | 99.40.1 |
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
| Pretrain | 96.20.2 | 41.90.8 | 95.90.1 | 0.90.3 | 24.21.0 | 40.62.1 | - | |
| Retrain | 96.50.0 | 31.30.3 | 95.90.0 | 1.60.4 | 27.00.4 | 34.40.9 | - | |
| Retrain | 96.30.1 | 39.70.7 | 95.80.0 | 2.21.2 | 23.90.8 | 41.31.6 | - | |
| L1-sparse (Jia et al., 2023) | 95.70.1 | 29.03.1 | 95.40.1 | 1.50.6 | 28.50.6 | 30.42.9 | 95.30.7 | |
| SalUn (Fan et al., 2024b) | 96.20.1 | 29.37.7 | 95.80.1 | 0.70.2 | 29.11.5 | 30.67.5 | 95.32.5 | |
| SCRUB (Kurmanji et al., 2024) | 96.50.2 | 35.11.7 | 95.90.0 | 0.60.2 | 26.61.1 | 35.92.0 | 97.50.4 | |
| MIU | 96.40.1 | 36.30.9 | 95.90.1 | 1.00.3 | 26.10.8 | 36.32.0 | 98.00.4 | |
| L1-sparse (Jia et al., 2023) | 95.60.0 | 37.33.8 | 95.40.1 | 0.70.3 | 26.70.9 | 34.84.8 | 97.41.4 | |
| SalUn (Fan et al., 2024b) | 96.10.1 | 42.910.5 | 95.80.0 | 0.60.5 | 23.92.8 | 41.19.8 | 96.12.0 | |
| SCRUB (Kurmanji et al., 2024) | 96.40.2 | 43.50.1 | 96.00.1 | 0.70.1 | 23.70.7 | 43.00.9 | 98.60.3 | |
| MIU | 96.30.2 | 43.20.6 | 96.00.0 | 1.20.1 | 24.00.5 | 41.31.1 | 99.00.2 |
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
| Pretrain | 96.40.1 | 44.06.6 | 95.90.1 | 0.70.1 | 23.51.3 | 41.15.2 | - | |
| Retrain | 96.70.0 | 20.21.0 | 95.90.0 | 7.21.4 | 31.50.6 | 25.41.9 | - | |
| Retrain | 96.40.1 | 33.61.1 | 95.70.1 | 4.31.3 | 25.60.7 | 35.91.9 | - | |
| L1-sparse (Jia et al., 2023) | 95.90.1 | 21.66.6 | 95.40.1 | 3.91.0 | 31.51.6 | 23.97.1 | 94.72.6 | |
| SalUn (Fan et al., 2024b) | 96.60.1 | 20.80.8 | 95.90.0 | 2.50.8 | 31.50.6 | 25.02.4 | 94.60.8 | |
| SCRUB (Kurmanji et al., 2024) | 96.70.1 | 27.21.2 | 95.90.0 | 3.10.5 | 31.30.4 | 26.30.7 | 96.00.6 | |
| MIU | 96.50.1 | 32.60.9 | 95.80.1 | 1.50.2 | 27.90.3 | 30.91.1 | 98.10.7 | |
| L1-sparse (Jia et al., 2023) | 95.90.0 | 33.22.9 | 95.20.1 | 2.60.7 | 28.30.5 | 33.11.8 | 98.40.6 | |
| SalUn (Fan et al., 2024b) | 96.40.1 | 38.16.7 | 95.80.2 | 1.91.2 | 25.91.1 | 36.97.2 | 97.10.4 | |
| SCRUB (Kurmanji et al., 2024) | 96.70.1 | 41.91.5 | 95.90.1 | 1.70.1 | 24.90.2 | 39.30.9 | 97.40.2 | |
| MIU | 96.70.1 | 35.22.0 | 95.80.0 | 4.21.1 | 26.91.4 | 34.32.7 | 98.80.4 |
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |